OpenStudy (anonymous):

The company discovered that it costs $28 to produce 2 camera cases, $50 to produce 4 camera cases, and $140 to produce 10 camera cases. Using the quadratic function, find the cost of producing 12 cameras.

OpenStudy (anonymous):

P = ax^2+bx+c, P is total production cost, x is number of cases. solve for a,b,c using the three values you are given,

OpenStudy (anonymous):

so, 28 = 4a+2b+c 50 = 16a+4b+c 140 = 100a+10b+c

OpenStudy (anonymous):

solve for a,b,c

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

do you know how to do matrices?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay 28 = 4a+2b+c ---------------------------------------->Eqaution 1 50 = 16a+4b+c ---------------------------------------> Eqn 2 140 = 100a+10b+c----------------------------------> eqn 3 do eqn 3 - eqn 1 and do eqn 2 - eqn 1 and post what you got here

OpenStudy (anonymous):

still don't know how to figure it out

OpenStudy (anonymous):

Can you subtract equation 1 from equation 2 and post it here?

OpenStudy (anonymous):

38=12a+2b

OpenStudy (anonymous):

how did you get 38?

OpenStudy (anonymous):

32*

OpenStudy (anonymous):

i mean 22*

OpenStudy (anonymous):

ok so repost the correct equation again

OpenStudy (anonymous):

22=12a+2b

OpenStudy (anonymous):

okay good. Now, the same way, subtract equation 1 from equation 3

OpenStudy (anonymous):

112=96a+8b

OpenStudy (anonymous):

okay. Now, can you solve the set of equations 12a + 2b = 22 96a + 8b = 112 ?

OpenStudy (anonymous):

can you help me with it.

OpenStudy (anonymous):

okay, have you learned how to solve two equations for two unknown variables in school?

OpenStudy (anonymous):

not that i can remember.

OpenStudy (anonymous):

you haven't done basic algebra in school yet?

OpenStudy (anonymous):

alright anyway, 12a + 2b = 22 ------------>eqnA 96a + 8b = 112----------->eqnB

OpenStudy (anonymous):

do you know substitution to solve for a and b?

OpenStudy (anonymous):

okay just do 4 times eqn A and post it here.

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

If you multiply both sides of equation A by 4, what do you get?

OpenStudy (anonymous):

48a+2b=88

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

that's multiplying both sides by 4.

OpenStudy (anonymous):

so 2b*4 = 2b?

OpenStudy (anonymous):

usually when i've done that, i've never done it to all. be more specific.

OpenStudy (anonymous):

48a+8b=88

OpenStudy (anonymous):

if a =4, then 4a =16 so we multiply both sides by 4. so if a+b = 4 then 4(a+b) =16

OpenStudy (anonymous):

okay good, so 48a + 8b = 88--------------->eqn. C Do you notice what we did there? we multiplied equation B by 4 so that both equation A and equation C have the term 8B in them. So now subtract equation C from equation A.

OpenStudy (anonymous):

i'm still completely confused

OpenStudy (anonymous):

what are you confused about?

OpenStudy (anonymous):

OK join me http://www.dabbleboard.com/draw?b=Guest654407&i=0&c=ce1b368e8fb4c87408d1e26d13c93f5bf92f9032 will explain in detail there

OpenStudy (anonymous):

okay i'm on there